OpenStudy (joehoffy):

Multiple choice The equation shows the relationship between x and y: y = −4x + 9 What is the slope of the equation? −9 −4 4 9

OpenStudy (fortytherapper):

y = mx + b Do you know what the m and b stand for?

OpenStudy (snuggielad):

B is your y intersect and M is your slope.

OpenStudy (snuggielad):

So, y = mx + b What's your slope?

OpenStudy (joehoffy):

lol nope i have no idea what ur talking about.

OpenStudy (snuggielad):

The equation you are using here is Y=mx+b Y and X are your variables M and B are there in place of numbers. So in this situation- \[y = −4x + 9 \] What is in the place of M?

jabez177 (jabez177):

\[y=mx+b\] \[y=-4x+9\]

jabez177 (jabez177):

What is M if you compare the 2 expressions I put up? :)

OpenStudy (snuggielad):

Okay, let's make it easier. \[y = ~mx + b\] \[y = -4x +9\]

OpenStudy (snuggielad):

What's in the place of M?

OpenStudy (snuggielad):

We are not going to give you the answer. It's against the code of conduct and punishable by banning. C'mon you've got to work with us here.

jabez177 (jabez177):

@Joehoffy?

OpenStudy (joehoffy):

aorry i was away XD

OpenStudy (joehoffy):

sorry*

OpenStudy (joehoffy):

personally, i think it is -4

OpenStudy (snuggielad):

And why do you think that?

OpenStudy (joehoffy):

i have a gut feeling about it.

jabez177 (jabez177):

You are correct because -4 is in the place of M in the y=Mx+b formula. :)

OpenStudy (snuggielad):

Your gut's not going to give you anything other than diarrhea.

jabez177 (jabez177):

Do you understand the question now? :)

jabez177 (jabez177):

If you would like I can give you are link to learn more about this. :-)

OpenStudy (joehoffy):

Thank you @jabez177 for teaching and not being rude about whatever i say :)

jabez177 (jabez177):

You're welcome. :) @SnuggieLad also helped. I just rounded things up. :)

myininaya (myininaya):

I think he was just trying to see if you understand it. I don't see it as rude if someone is trying to teach you and assist you.

jabez177 (jabez177):

The "Guts and Diarrhea" part. :p

myininaya (myininaya):

I think that part was funny... Guts and diarrhea do have a high correlation together.

myininaya (myininaya):

:p

OpenStudy (just_one_last_goodbye):

Nicely put @myininaya and @jabez177 lets just leave things alone and move on with the day ^_^ come on.... the day is almost over :/ :) lets just move on right?

myininaya (myininaya):

Humor is great when it comes to learning. :)

jabez177 (jabez177):

^ True. It makes learning fun! :D

OpenStudy (just_one_last_goodbye):

Learning is lacking humor now and days... thats why there is so much violence in public schools and bullying just cause kids want to find "entertainment"